OpenStudy (anonymous):

solve the system of equations by graphing. -1/3x + y = -1 y =4 + 1/3x

OpenStudy (anonymous):

u have to plot this y =4 + 1/3x?

OpenStudy (anonymous):

-1/3x + y = -1

OpenStudy (anonymous):

oh got it...there r two equations

OpenStudy (anonymous):

y =4 + 1/3x y = -1+3/x

OpenStudy (anonymous):

u know to plot y=1/x?

OpenStudy (anonymous):